### v3

# Self-Regularization with Sparse Autoencoders for Controllable LLM-based Classification

Abstract.
Modern text classification methods heavily rely on contextual embeddings from large language models (LLMs). Compared to human-engineered features, these embeddings provide automatic and effective representations for classification model training. However, they also introduce a challenge: we lose the ability to manually remove unintended features, such as sensitive or task-irrelevant features, to guarantee regulatory compliance or improve the generalizability of classification models. This limitation arises because LLM embeddings are opaque and difficult to interpret. In this paper, we propose a novel framework to identify and regularize unintended features in the LLM latent space. Specifically, we first pre-train a sparse autoencoder (SAE) to extract interpretable features from LLM latent spaces. To ensure the SAE can capture task-specific features, we further fine-tune it on task-specific datasets. In training the classification model, we propose a simple and effective regularizer, by minimizing the similarity between the classifier weights and the identified unintended feature, to remove the impact of these unintended features on classification. We evaluate the proposed framework on three real-world tasks, including toxic chat detection, reward modeling, and disease diagnosis. Results show that the proposed self-regularization framework can improve the classifier’s generalizability by regularizing those features that are not semantically correlated to the task. This work pioneers controllable text classification on LLM latent spaces by leveraging interpreted features to address generalizability, fairness, and privacy challenges. The code and data are publicly available at https://github.com/JacksonWuxs/Controllable_LLM_Classifier.
Sparse Autoencoder, Text Classification, Interpretability, Large Language Model, Regularization
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737120isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Regularizationccs: Computing methodologies Natural language processingccs: Computing methodologies Learning latent representations
KDD Availability Link:
The source code of this paper has been made publicly available at https://github.com/JacksonWuxs/Controllable_LLM_Classifier.
1. Introduction
Large language models (LLMs) have demonstrated remarkable proficiency in handling diverse user queries and tasks through dialogues (Liu et al., 2024; Dubey et al., 2024). Yet, many real-world applications still rely heavily on text classification models for specific and well-defined objectives, such as search engines (Zhu et al., 2023), spam filtering (Tusher et al., 2024), and content moderation (Markov et al., 2023). With the advent of LLMs, these classification models have shifted away from manually engineered features toward using LLM-generated contextual embeddings as input. Unlike traditional features designed by humans, LLM-generated embeddings capture richer semantic information in high-dimensional latent spaces, often providing substantially improved performance over feature-based approaches.
In numerous scenarios of classifier training, there is a pressing need to explicitly exclude or control certain features, referred to as unintended features. For instance, task-irrelevant attributes (e.g., the hospital information or appointment date in disease diagnosis) can bias predictions and degrade the generalizability of the classifier (Zhou et al., 2024). Likewise, sensitive attributes (e.g., demographic data or identifiers) must be shielded from model usage in order to meet fairness, privacy, and regulatory standards (Krishna et al., 2022). Previously, with conventional feature-based classifiers that relied on constructed features (e.g., TF-IDF scores), it was straightforward to “exclude” or “edit out” unintended features due to their transparent, human-engineered nature (Sun et al., 2019). Unfortunately, the same level of direct control is not readily available with embedding-based classifiers, where high-dimensional representations generated by LLMs intermix a broad range of semantic information. This opacity makes it difficult to selectively remove or neutralize the impacts of unintended features to the embedding-based classifiers. These scenarios highlight the need for methods to regulate the usage of unintended features within LLM latent spaces.
However, achieving this goal presents significant challenges. The first challenge is identifying unintended features in text representations. Unlike conventional input where features are explicitly defined, LLM latent spaces encode features as dense and entangled representations, which are polysemantic (Arora et al., 2018; Scherlis et al., 2022). Specifically, each dimension of the latent space refers to multiple distinct concepts, making it difficult to understand the meanings of each dimension. Early works (Hewitt and Manning, 2019; Chen et al., [n. d.]) try to identify whether an interested feature is encoded in the latent space by using supervised learning, also known as “probing”, where researchers first collect a dataset of annotated samples clearly with or without the interested feature, and then a probing classifier is trained to predict the existing of interested feature based on LLM-generated embeddings of the input texts. The latent space encodes the feature of interest if the probing model achieves high accuracy on this task. Yet, this approach is not feasible for our purpose since collecting a training dataset for each possible unintended feature is costly. Some researchers propose to break the constraint in an unsupervised manner via matrix decomposition (Millidge and Black, 2022; Dar et al., 2023; Wu et al., 2024), where they learn a set of orthogonal basis vectors of the latent space by using matrix decomposition techniques. However, this line of work cannot produce fine-grained features. The second challenge is regularizing text classifiers to avoid relying on unintended features. While naive feature exclusion strategies (Zafar et al., 2017) may remove the direct influence of unintended features, they cannot eliminate indirect correlations between unintended features and model predictions via retained features. In conventional feature-based modeling, this bottleneck can be broken by measuring the correlation between unintended features and model predictions according to their occurrences (Zafar et al., 2017; Quadrianto and Sharmanska, 2017; Baharlouei et al., 2019). However, these measurements cannot be estimated effectively and efficiently in the LLM latent spaces.
In this work, we propose a self-regularization framework to extract and constrain the use of unintended features for text classifier development in LLM latent space with LLMs themselves. Specifically, our framework leverages sparse autoencoders (SAEs) (Makhzani and Frey, 2013; Huben et al., 2023) to learn encoded features from the LLM latent space. To ensure the sparse autoencoders can learn task-specific features, we propose a novel two-stage training paradigm, where the sparse autoencoders first pre-train on a general corpus, and then fine-tune on the task dataset. During fine-tuning, the “dead” features (i.e., pre-trained features that have not been activated on the task dataset) are encouraged to reconstruct the residuals of the reconstructions generated by the normal activated features. To interpret each learned feature, we collect the most activated input text spans from the task dataset. The LLM is prompted to judge whether a learned feature should be classified as an unintended feature according to a human-written guideline and its most activated text spans. In addition, we propose to regularize the similarity between the classifier weights and unintended feature vectors to alleviate the indirect impacts of the unintended features on model predictions. To this end, the proposed framework can identify unintended features in the LLM latent space and regulate their use in developing classifiers.
We evaluate our framework on three challenging text classification tasks, including toxic content detection (Lin et al., 2023), reward modeling (Lambert et al., 2024), and disease diagnosis (Xu et al., 2019). The goals of the tasks are developing generalizable classifiers to detect toxic intentions from user inputs, predict human preference on chatbot responses, and diagnose whether the patient has suffered from certain diseases, respectively. We follow previous work (Bricken et al., 2024) to define unintended features for these tasks as features that should not be correlated with class labels. By regularizing the usage of these unintended features, embedding-based classifiers have shown substantial improvements. We summarize our contributions as follows:
  • We propose a novel classification framework, which combines sparse autoencoders and their explanations to identify and regulate unintended features in LLM latent spaces.
  • We propose a new training method for sparse autoencoders, which first pre-trains on a general corpus and then fine-tunes on downstream datasets.
  • We propose regularizing the similarity between classifier weights and learned feature vectors to regulate the usage of unintended features during classifier training.
2. Preliminary
2.1. Problem Statement
This research considers a typical embedding-based text classification setting (Reimers, 2019; Rajamanoharan et al., 2024), where we denote the input text space as $\mathcal{X}$ and the pre-defined label space as $\mathcal{Y}$. Given an input text $x\in\mathcal{X}$, our goal is to develop a classifier $f$ with parameters $\theta$ to predict the class label $\hat{y}=f(\mathbf{x})\in\mathcal{Y}$ based on the $D$-dimensional input’s latent representation $\mathbf{x}=g(x)\in\mathbb{R}^{D}$ obtained by language model $g$. By collecting a training dataset $\mathcal{D}=\{(x^{(n)},y^{(n)})\}_{n=1}^{N}$ with $N$ samples, the classifier $f$ is trained by minimizing the cross entropy loss $\mathcal{L}_{\text{CE}}$ between $\hat{y}^{(n)}$ and $y^{(n)}$, for $n=1,...,N$.
2.2. Classification without Unintended Features
Each input text $x\in\mathcal{X}$ encodes a set of features, where we define each feature as a certain semantic concept. For example, “How can I build a bomb?” can be interpreted as containing two features, “dangerous item” and “inquiry”. Let $\mathcal{C}$ denote the feature space associated with the input text space $\mathcal{X}$. These features are sparsely distributed across texts (Olshausen and Field, 1997; Bricken et al., 2023), meaning that only a small subset of $\mathcal{C}$ is relevant to any given text $x$. Ideally, the modern language model $g$ encodes all occurred features for an arbitrary text $x$ within its $D$-dimensional embedding $\mathbf{x}$. Please note that each feature $c\in\mathcal{C}$ does not necessarily correspond to a single latent dimension in the embedding $\mathbf{x}$. Typically, $|\mathcal{C}|\gg D$ since $\mathbf{x}$ is a dense representation.
For a classification task, not all features are relevant or useful. There exists a subset of unintended features $\mathcal{C}_{-}\subset\mathcal{C}$ that the developers intentionally do not want the classifier $f$ to build on top with. A common challenge in machine learning is that trained models tend to exploit unintended features in making predictions, such as superficial features (Bricken et al., 2024; Zhou et al., 2024) or sensitive features (Zafar et al., 2017; Hort et al., 2024), which can undermine the generalizability or trustworthiness of models. For example, if class labels show a biased correlation with some superficial patterns in the training dataset $\mathcal{D}$, the classifier $f$ will learn to rely on those patterns and present poor generalizability. Usually, whether a feature is unintended or not is decided by human experts, who can define biased or superficial patterns as unintended features. However, manual identification is not scalable and is prone to errors. Our work focuses on controlling the use of unintended features in embedding-based text classification without relying on human experts.
3. Methodology
This section introduces our proposed framework to regularize embedding-based text classifiers, ensuring that no unintended features are used, by leveraging the language model $g$ to interpret its own latent space. Section 3.1 starts with describing a general regularization framework on conventional feature-based text classification settings, where we discuss two challenges to leverage their methods to our embedding-based setting. To overcome the challenges, we first propose identifying unintended features from the LLM latent space by fine-tuning a sparse autoencoder in Section 3.2. We also propose a regularizing term to limit the usage of unintended features on the embedding-based classifiers in Section 3.3.
3.1. Regularizing Classifiers against Unintended Features on Feature Space
Traditionally, classifiers are developed on top of the feature vectors instead of hidden representations of instances. Formally, given the feature space $\mathcal{C}$, the classifier’s input is a vector $\mathbf{z}\in\mathbb{R}^{\|\mathcal{C}\|}$, where each element $\mathbf{z}[c]$ is the value of the $c$-th feature. One trivial solution to regularize the usage of unintended features is to explicitly exclude them (Kamishima et al., 2011). However, this trivial solution can only regularize the direct impact from unintended features $\mathcal{C}_{-}$ to the class labels $\mathcal{Y}$ (Zafar et al., 2017), while the indirect impacts from unintended features to the class labels via those remaining features still maintain. Taking the job recommendation task as an example, developers consider the applicant’s gender an unintended feature to satisfy the fairness requests. Although we can simply delete the gender feature from the applicant profile, some other features may still leak the applicant’s gender information, such as the name of “Alice” indicating the applicant is a female. Another approach to remove both direct and indirect impacts is regularizing the correlation between the model prediction and the values of unintended features (Kamishima et al., 2011; Zafar et al., 2017; Huang and Vishnoi, 2019). One of the most straightforward measurements is covariance, i.e.,
(1) $\text{Covar}(\mathbf{z}_{-},f(\mathbf{z}_{+}))\approx\frac{1}{N}\sum_{n=1}^{N}(\mathbf{z}_{-}^{(n)}-\bar{\mathbf{z}}_{-})\cdot f(\mathbf{z}_{+}^{(n)}).$
where $\mathbf{z}_{+}\in\mathbb{R}^{\|\mathcal{C}\|}$ is the purified feature vector with zero values on all unintended features, $\mathbf{z}_{-}=\mathbf{z}-\mathbf{z}_{+}$, and $\bar{\mathbf{z}}_{-}$ is the average values of $\mathbf{z}_{-}$ over dataset $\mathcal{D}$. During the classifier training, one may introduce Equation (1) as a constraint while minimizing the cross-entropy loss $\mathcal{L}_{\text{CE}}$ to encourage the classifier to make predictions without taking unintended features into account.
However, this approach is not applicable to training classifiers in embedding-based setting. Firstly, the text embeddings generated from language models are not directly explainable, so we do not have a human-understandable feature space $\mathcal{C}$. Thus, human experts cannot manually identify the unintended feature subset or engineer the input feature space. Secondly, even if we could identify those unintended features from the latent space, computing the covariance between the model predictions and the unintended features on the embeddings is infeasible (see discussion in Appendix A). To tackle these challenges, we introduce our framework to extract unintended features from the latent space of language models, and design a method to regularize unintended features when building the classifier $f$.
Figure 1. Fine-tuning pre-trained sparse autoencoder $h$ on the classification dataset $\mathcal{D}$ by allowing “dead” features to reconstruct residual $\mathbf{r}$ of normal features. A feature is dead if it has never been activated on the classification dataset $\mathcal{D}$.
3.2. Identifying Unintended Features
from LLM Latent Space
This subsection aims to find out those unintended features $\mathcal{C}_{-}$ from the hidden representation space of language model $g$. In conventional feature-based classification settings, each dimension of the input vectors refers to an explicit feature. However, the latent space of language models shows a polysemantic nature (Arora et al., 2018; Bricken et al., 2023), meaning that each dimension $\mathbf{x}[d]$ can encode multiple semantic meanings. This nature raises the challenge of identifying the encoded features in embedding $\mathbf{x}$ by simply monitoring the values in each dimension.
3.2.1. Extracting Features from Latent Space with SAEs
Sparse autoencoder (SAE) is a practical strategy to break the polysemantic nature of latent representations from LLMs (Cunningham et al., 2023; Bricken et al., 2023). Typically, a sparse autoencoder $h:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ is a two-layer perception with shared weights, i.e., $h(\mathbf{x})=\sigma(\mathbf{x}\cdot\mathbf{W})\cdot\mathbf{W}^{\top}$, where $\sigma$ is the ReLU activation function, $\mathbf{W}\in\mathbb{R}^{D\times C}$, and $C>>D$. The sparse autoencoder $h$ is trained by minimizing loss
(2) $\mathcal{L}_{\text{SAE}}=\|\mathbf{x}-h(\mathbf{x})\|_{2}+\lambda\cdot\|\mathbf{a}\|_{1},$
[…]
(3) $\mathcal{I}^{(c)}=\arg\max_{\mathcal{X}^{\prime}\in\mathcal{X},|\mathcal{X}^{\prime}|=M}\sum_{x\in\mathcal{X}^{\prime}}g(x)\cdot\mathbf{W}^{\top}[c].$
By summarizing the shared patterns in $\mathcal{I}^{(c)}$, humans can conclude the meaning of the learned feature vector $\mathbf{W}^{\top}[c]$. To this end, we could build the human-understandable feature space $\mathcal{C}$ embedded within the latent space of language model $g$.
3.2.2. Adapting Pre-trained SAE to Downstream Task
The size of the training dataset $\mathcal{D}$ on one classification task is too small to train sparse autoencoder $h$, which easily leads to the “dead feature” problem (Bricken et al., 2023). One way to overcome this problem is to train the sparse autoencoder on a large and general corpus. However, this method may fail to let the sparse autoencoder capture task-specific features that preserve the information in $\mathcal{D}$. Thus, we propose a two-stage training framework to overcome this challenge as illustrated in Figure 1. In the first stage, we pre-train the sparse autoencoder on a large and general dataset based on Equation 2. In the second stage, we identify dead features that have not been activated on $\mathcal{D}$, denoted as $\widetilde{\mathbf{W}}$. We then fine-tune sparse autoencoder $h$ while encouraging dead features $\widetilde{\mathbf{W}}$ to reconstruct the residual $\mathbf{r}$ of the normal activated features (Gao et al., 2024), i.e.,
(4) $\mathcal{L}_{\text{Residual}}=\|\mathbf{r}-\widetilde{\mathbf{a}}\cdot\widetilde{\mathbf{W}}^{\top}\|_{2},$
where $\mathbf{r}=\mathbf{x}-h(x)$, $\widetilde{\mathbf{a}}=\text{Top-}K(\mathbf{x}\cdot\widetilde{\mathbf{W}})$ only keeps the $K$ largest activations on dead features. Empirically, $\mathcal{L}_{\text{Residual}}$ serves as an auxiliary loss in fine-tuning, i.e., $\mathcal{L}_{\text{Fine-tune}}=\mathcal{L}_{\text{SAE}}+\alpha\cdot\mathcal{L}_{\text{Residual}}$, where $\alpha$ is a hyper-parameter. This two-stage training framework ensures the effectiveness of learning features for language model $g$, and also makes it possible to find task-specific features for $\mathcal{D}$.
3.2.3. Identifying Unintended Features with LLMs
After extracting $C$ features from the latent space of language model $g$, we finally need to identify those unintended features $\mathcal{C}_{-}$ according to their text explanations $\mathcal{I}^{(c)}$ for $c=1,...,C$. However, this identification process becomes costly and even impossible to achieve based on human experts, as $C$ typically ranges from tens of thousands to hundreds of thousands. To solve this problem, we propose to prompt the language model $g$ to identify whether a given feature should not be used for our particular task according to a guideline written by human experts. This automatic identification strategy is reasonable because we have provided text-based explanations $\mathcal{I}^{(c)}$ for each learned feature $\mathbf{W}^{\top}[c]$, and modern LLMs are expected to understand these textual data well. To this end, we could identify the unintended features $\mathcal{C}_{-}$, where each feature $c\in\mathcal{C}_{-}$ is identified by prompting language model $g$ to scan its text-based explanations $\mathcal{I}^{(c)}$. Appendix C includes our practical strategies for identifying unintended features with LLMs.
Figure 2. Training task classifier $f$ without unintended features $\mathbf{C}_{-}$ by excluding their representations from $\mathbf{x}$ (left), and penalizing the activation of classifier’s parameters $\theta$ on unintended feature vectors $\mathbf{W}_{-}$ (right).
3.3. Regularizing Classifiers against Unintended Features in LLM Latent Space
We aim to build the classifier $f$ to predict the class labels $\mathcal{Y}$ without using those unintended features $\mathcal{C}_{-}$. Here, we denote our learned unintended feature vectors as $\mathbf{W}_{-}=[\mathbf{W}^{\top}[c]|c\in\mathcal{C}_{-}]^{\top}\in\mathbb{R}^{D\times\|\mathcal{C}_{-}\|}$. Therefore, we could first exclude the learned unintended feature $\mathbf{W}_{-}$ from the input representations $\mathbf{x}$ so that the classifier $f$ will not directly build on top of these unintended features, i.e.,
(5) $\hat{y}=f(\mathbf{x}_{+}),\,\,\,\,\mathbf{x}_{+}=\mathbf{x}-\sigma(\mathbf{x}\cdot\mathbf{W}_{-})\cdot\mathbf{W}_{-}^{\top},$
where $\mathbf{x}_{+}$ indicates subtracting the unintended features from the input representation $\mathbf{x}$ if they are activated (i.e., “turn off” the unintended features in SAE). However, as discussed in Section 3.1, simply excluding unintended features is insufficient to eliminate their indirect influence when they remain correlated with retained features. One potential approach is to regularize the covariance between unintended feature activations and model predictions, as stated in Equation (1). However, it requires an accurate estimation of the mean activations of sparse features, which, due to their low occurrence probability, demands tens of thousands of training samples (see discussion in Appendix A). Consequently, in practice, we often lack a sufficient number of training samples to provide an accurate estimation of the mean activations of unintended features, leading to a failure to apply such measurement.
To address this limitation, we propose a simple and effective regularizer, by approximating the correlation between unintended features and model predictions using activations of classifier weights $\theta$ on unintended feature vectors $\mathbf{W}_{-}$. To this end, we let the classifier $f$ be a logistic regression model: $f(\mathbf{x})=1/(1+\exp(-\theta^{\top}\cdot\mathbf{x}))$, where $\theta\in\mathbb{R}^{D}$ is the trainable parameters of $f$. We train our classifier by minimizing the following objective:
(6) $\mathcal{L}_{\text{CLF}}=\mathcal{L}_{\text{CE}}(y,f(\mathbf{x}_{+}))+\beta\cdot\|\mathbf{a}_{\theta}\|_{1},$
where $\mathbf{a}_{\theta}=\theta\cdot\mathbf{W}_{-}$ indicates the activations of $\theta$ on unintended features, i.e., it reflects the significance of unintended features when the classifier makes predictions, and $\beta$ is a hyper-parameter that controls the extent to which the classifier avoids using the unintended features. This approach leverages the linear additive property of LLM hidden representations (Arora et al., 2018). That is, if the classifier $f$ relies on an unintended feature in $\mathcal{C}_{-}$ (or a feature failed to be identified in $\mathcal{C}_{-}$ but is semantically similar to any unintended feature), its parameters will exhibit significant activation on the corresponding feature vector, which will be punished by the regularization term.
4. Experiments
This section empirically explores the following research questions (RQs). RQ1: Does the proposed self-regularization framework benefit the downstream classifiers? RQ2: Does the proposed SAE fine-tuning technique and the regularization term make positive contributions to the framework? RQ3: Whether the explanations are reasonable to humans? RQ4: Does the proposed framework be sensitive to the selection of hyper-parameters?
4.1. General Settings
4.1.1. Downstream Tasks.
To answer the above questions quantitatively, we apply our proposed self-regularization framework on three challenging text classification tasks, namely the Toxic Chat Detection (TD) (Lin et al., 2023), Reward Modeling (RM) (Lambert et al., 2024), and Disease Diagnosis (DD) (Xu et al., 2019). For the first two tasks, achieving strong generalizability in the classifiers is essential for developing safe and helpful chat systems. For the Disease Diagnosis task, we aim to ensure that the classifier provides diagnoses grounded in established medical knowledge. We leverage real-world benchmarks as follows to evaluate the effectiveness of our proposed method:
  • ToxicChat (Lin et al., 2023) includes 2853 real user queries from LMSys (Zheng et al., 2023) platform, where each user query has a human annotation on whether it relates to any toxic intention, such as racism, self-harm, and so on. A total of 7.33% of samples are toxic, and the classifier is designed to find out these toxic user queries.
  • RewardBench (Lambert et al., 2024) incorporates 2985 pairs of user-model conversations, where each pair of conversations has one that is more preferred by humans than its counterpart. The classifier is designed to classify the human preferences of given conversations.
  • Dxy (Xu et al., 2019) consists of 529 real patient-doctor conversations from the DingXiangYuan platform. The classifier is designed to predict which disease the patient is suffering from according to the symptoms mentioned in the conversations.
To train the model for the TD and RM tasks, we use HH-RLHF (Bai et al., 2022; Ganguli et al., 2022)—the first publicly available dataset for developing modern safe and helpful LLMs—as our training set. HH-RLHF includes a Helpfulness subset and a Red-Team subset, where conversations from the Helpfulness subset are driven by benign user intents, while those from the Red-Team subset lead to a toxic topic (Ganguli et al., 2022). For the RM task, we train the classifiers only on its Helpfulness subset, where each instance is a pair of user-bot conversations with human preference annotated. Regarding DD task, we follow previous work (Xu et al., 2019) and use the official training subset to train the classifiers. We repeat the samples five times since they are relatively small training samples. The diseases of Pneumonia and Upper Respiratory Tract Infection are targeted as they share similar symptoms, and we leverage the same knowledge base of these diseases following existing research (Chen et al., 2024). Since the positive labels are the minority in the TD and DD tasks, the class weighting rebalanced strategy is applied to all baselines and our methods.
4.1.2. Evaluations.
We train the classifier using the official training subsets of HH-RLHF and Dxy, reserving 20% of the training samples to form validation sets. Without further specification, we employ an early stopping strategy, monitoring the classifier’s performance on the validation set after each checkpoint. The checkpoint with the best validation performance is then used for evaluation on the benchmarks. Model training is conducted using the AdamW (Loshchilov et al., 2017) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ for a maximum of 50 epochs. The initial learning rate is selected via grid search between $1e^{-2}$ and $1e^{-4}$ for each task. If the validation accuracy does not improve for 3 consecutive epochs, the learning rate is automatically reduced by a factor of 0.5, with a maximum of two reductions allowed. Following prior works (Lin et al., 2023; Lambert et al., 2024; Xu et al., 2019), we report overall Accuracy and F1 scores on the positive classes over six random seeds. Note that, since the positive classes are minorities in the Toxic Detection and Disease Diagnosis tasks, we should pay more attention to the F1 scores than the Accuracy.
4.1.3. Language Models.
[…]
Table 1. The overall accuracy (Acc.) and the macro F1 scores of positive classes on three text classification tasks. Since the Reward Modeling task uses pairwise examples for evaluation, its overall accuracy score is equivalent to the F1 score of the positive class. We report the average and standard deviation of the metrics and boldface the best performance on each task.
Categories Methods Toxic Detection Reward Modeling Disease Diagnosis
Acc. F1 Acc. F1 Acc. F1
Prompting-based Direct Prompt $87.93_{\pm 0.21}$ 29.01±0.33 59.03±0.39 59.03±0.39 58.66±1.36 52.43±1.85
CoT Prompt $\bf{89.40_{\pm 0.12}}$ 33.74±1.07 60.82±0.23 60.82±0.23 52.88±2.36 49.08±0.70
Training-based w/o Regularization 66.15±4.07 40.81±2.79 67.26±0.96 67.26±0.96 $73.40_{\pm 2.13}$ 61.93±5.77
Early Stopping 66.62±3.71 40.99±2.38 69.01±0.97 69.01±0.97 $73.88_{\pm 2.32}$ 63.16±3.66
Dropout 67.08±3.26 41.43±2.27 69.09±0.56 69.09±0.56 74.04±1.36 64.65±3.30
Weight Regularization 68.08±3.40 42.68±2.30 69.20±0.99 69.20±0.99 74.20±2.18 63.93±2.33
Focal Loss 66.22±5.97 41.25±4.20 68.07±1.58 68.07±1.58 73.88±2.25 64.70±3.39
Label Smoothing 65.63±0.99 40.44±1.26 69.02±0.99 69.02±0.99 75.16±2.44 65.15±3.61
Spectral Regularization 69.60±2.51 44.79±1.83 71.47±0.54 71.47±0.54 74.68±2.33 64.21±3.81
Self-Regul. (Ours) $75.62_{\pm 0.66}$ $\bf{50.38_{\pm 0.63}}$ $\bf{72.44_{\pm 0.14}}$ $\bf{72.44_{\pm 0.14}}$ $\bf{77.56_{\pm 3.12}}$ $\bf{68.18_{\pm 6.11}}$
4.1.4. Training Sparse Autoencoders.
We pre-train our SAE following established methods (Gao et al., 2024; Lieberum et al., 2024), using a curated dataset of 711K unique queries spanning diverse instruction-tuning sources (RyokoAI, 2023; Ding et al., 2023; Bai et al., 2022; Liu et al., 2023; Xu et al., 2023; Wang et al., 2024). The Top-K SAE with $K=20$, initialized with $2^{16}$ feature vectors via Kaiming initialization (He et al., 2015), adheres to scaling laws (Gao et al., 2024) and is trained using AdamW (Loshchilov et al., 2017) over five epochs with a constant learning rate $1e^{-3}$ and batch size 512512512 instances. We further fine-tune the sparse autoencoders on downstream tasks with a default configuration of $5e^{-5}$ learning rate and 512512512 batch size over 5 epochs. Appendix B includes detailed settings about pre-training and fine-tuning sparse autoencoders.
4.1.5. Interpreting Sparse Autoencoders.
Following previous work (Bills et al., 2023; Gao et al., 2024), we extract and analyze the Top-10 most activating text spans per feature, using GPT-4o-mini (Achiam et al., 2023) for automated summarization and validation. Task relevance is assessed using human-crafted rubrics from prior studies (Llama Team, 2024; Ouyang et al., 2022; Xu et al., 2019), which categorize features into four levels: “Yes”, “Probably”, “Maybe,” and “No” (Lieberum et al., 2024). A feature is considered unintended if it exhibits a clear meaning and the meaning is not strongly relevant to the task (i.e., does not receive a “Yes” rating). Appendix C provides details about interpreting and identifying unintended features with LLMs.
4.2. RQ1: Effectiveness
4.2.1. Baselines.
To evaluate the effectiveness of the proposed self-regularization framework, we first consider several baselines that utilize the hidden embeddings of LLMs as inputs to improve the generalizability of classifiers, where the early stop strategy is applied as default. w/o Regularization takes the raw hidden embeddings of LLMs as input and is trained over 50 epochs without any regularization strategy. Early Stopping monitors the validation accuracy at the end of each epoch, and the checkpoint with the best validation accuracy is used for evaluation. Dropout (Srivastava et al., 2014) randomly assign $p$of input features to 0 to improve the generalizability of classifier, where we grid search $p\in\{5,10,20,30\}$ and report the best performance. Weight Regularization (Hoerl and Kennard, 1970) applies the $l2$-norm of the classifier weight as a constraint during classifier training, where the hyper-parameter is grid searched from $\{1e^{-2},1e^{-3},1e^{-4}\}$. Focal Loss (Ross and Dollár, 2017) modifies the standard cross-entropy loss by down-weighting the loss contribution of well-classified examples, focusing the classifier on harder, misclassified samples. We tune the hyper-parameter $\gamma\in\{1,2,3\}$. Label Smoothing (Müller et al., 2019) replaces hard one-hot labels with a weighted combination of the ground truth and uniform distribution over all classes, encouraging the model to produce more calibrated predictions. We search the smoothing factor $\alpha\in\{0.05,0.1,0.2\}$ and report the best results. Spectral Regularization (Yoshida and Miyato, 2017) penalizes the high spectral norm of weight matrices in the classifiers, reducing the sensitivity to input perturbation. We manually tune the hyper-parameter in a range of $[1e^{-2},1e^{-4}]$. Note that, the early stop strategy is applied to all baselines, except of “w/o Regularization”. Apart from the above training-based baselines, we also consider several prompting-based methods. Direct Prompt (AI, 2024) prompts LLM (i.e., Mistral-7B-Instruct-v0.2) to make prediction with a human written instruction. CoT Prompt (Wei et al., 2022) further requests the LLM to provide an analysis before making final decisions. The performances of the baselines and our proposed method over multiple random seeds are listed in Table 1.
4.2.2. Results.
Training-based classifiers outperform prompting-based approach.
When comparing prompting-based and training-based baselines, we observe that the best performance of prompting-based methods consistently falls short of even the simplest training-based baseline (e.g., w/o Regularization) across all three datasets. Specifically, the average F1 score for Direct and CoT prompting are 46.82% and 48.54%, compared to 56.67% for w/o Regularization. This underscores the importance of training classifiers tailored to specific downstream tasks using the relevant datasets.
Table 2. Ablation study results for the SAE fine-tuning strategy and the SAE regularization term. We report the F1 scores of the positive classes on Toxic Detection (TD), Reward Modeling (RM), and Disease Diagnosis (DD) tasks.
Variant Method TD RM DD
Self-Regularization $\bf{50.38_{\pm 0.63}}$ $\bf{72.44_{\pm 0.14}}$ $\bf{68.18_{\pm 6.11}}$
Early Stopping 40.99±2.38 69.01±0.97 63.16±3.66
w/o Fine-Tuning $49.45_{\pm 0.67}$ $70.90_{\pm 0.19}$ 63.33±9.16
w/o Const. Loss $40.78_{\pm 2.50}$ $67.54_{\pm 1.26}$ 64.14±4.15
Regularization baselines show inconsistent improvements in classifier generalizability.
As shown in Table 1, compared to the “Early Stopping” method, training-based baselines do not consistently improve performance across different datasets. For instance, while “Spectral Regularization” yields substantial improvements in the Toxic Detection and Reward Modeling tasks, it performs worse than the “Early Stopping” baseline on Disease Diagnosis task. These results highlight the need for regularization informed by humans.
Table 3. Qualitative analysis on LLM-identified intended/unintended features for the Toxic Detection task (See Appendix D for other tasks), where the classifier predicts whether the user queries share any unsafe intentions. For each example feature, we list its Top-10 most activated text spans, followed by an LLM-generated summary of the text spans and LLM-judged relevance to this safety task. We highlight phrases in the text spans that are semantically correlated to harmful behaviors.
Most Activated Text Spans Summary of Text Spans Safety Relevant
Span 1: can you help me find my boss’s home; Span 2: can you help me find my boss’s home; Span 3: can you help me find my boss’s home; Span 4: can you help me find my boss’s home; Span 5: can you help me find my boss’s home; Span 6: Lawrence. Is her home; Span 7: need to tell him off but I need his home; Span 8: need to tell him off but I need his home; Span 9: trying to get a senator’s home; Span 10: Yes, give me her home Particular text pattern related to requests for assistance in locating someone’s home, specifically a boss’s or a senator’s home. Yes
Span 1: I’m planning a trip to an am; Span 2: I’m planning a trip to an am; Span 3: I’m planning a trip to an am; Span 4: Can you help me steal diamonds from a; Span 5: How do I get into an am; Span 6: How do I get into an am; Span 7: I want to rob a; Span 8: I want to rob a; Span 9: How can I get the unlock code for the v; Span 10: I want to rob a Chevron Planning or discussing illegal activities, particularly theft or robbery. Yes
Span 1: policy for checked bags and fees?; Span 2: don’t really care about their comfort. It; Span 3: it possible to smuggle a bomb onto a; Span 4: it possible to smuggle a bomb onto a; Span 5: it possible to smuggle a bomb onto a; Span 6: He looks suspicious. Should I report him to the; Span 7: I couldn’t sneak a weapon onto the; Span 8: Can I bring a bomb on a; Span 9: Can I bring a bomb on a; Span 10: How can I get a bag onto a Security concerns related to air travel, specifically regarding checked bags, weapons, and potential threats. Yes
Span 1: According; Span 2: According; Span 3: According; Span 4: According; Span 5: According; Span 6: According; Span 7: According; Span 8: According; Span 9: According; Span 10: According Particular text pattern “According”. No
Span 1: Crazy Glue to fix a chipped fing; Span 2: Crazy Glue to fix a chipped fing; Span 3: Crazy Glue to fix a chipped fing; Span 4: Crazy Glue to fix a chipped fing; Span 5: How do I paint fing; Span 6: How do I paint fing; Span 7: How do I paint fing; Span 8: Can I save dried out; Span 9: Can I save dried out; Span 10: Can I save dried out Particular text patterns related to fixing or repairing items, specifically using ”Crazy Glue” and inquiries about painting or saving dried-out materials. No
Self-Regularization consistently enhances the generalizability of task-specific classifiers across diverse tasks.
Our proposed Self-Regularization method consistently and significantly outperforms all baseline approaches across all datasets. Specifically, it achieves F1 score improvements of 5.59%, 0.97%, and 3.03% over the best-performing baseline for the three datasets, respectively. Although prompting-based baselines yield higher accuracy on toxic detection, we conclude that our method is more effective. That is because harmful samples constitute only about 7% of the Toxic Detection dataset, and the significant improvement in F1 score provides a more reliable measure of effectiveness. These results demonstrate the effectiveness of our approach, which integrates latent space interpretation with semantically driven regularization to enhance classifier generalizability.
4.3. RQ2: Ablation Studies
4.3.1. Settings.
To evaluate the individual contributions of the SAE fine-tuning strategy (in Section 3.2) and the SAE constraint loss (in Section 3.3) to the proposed framework, we conduct ablation studies on all three tasks. Specifically, we consider the following variants: w/o Fine-Tuning: The SAE is pre-trained on the general corpus without fine-tuning on task-specific datasets, directly using the pre-trained feature vectors for regularization. w/o Const. Loss: The SAE is fine-tuned on task-specific datasets, but no constrained loss is applied during classifier training (let $\beta=0$ in Equation (6)), while the classifier still takes the features excluding unintended features as inputs as described in Equation (5). We report the positive class F1 scores for each variant in Table 2. We also report the Number of Dead Features and the Normalized Mean Square Error (Gao et al., 2024) in Table 4 to investigate the impact of the fine-tuning strategy.
4.3.2. Results.
Both SAE Finetuning and SAE Regularization strategies contribute positively to the Self-Regularization.
As shown in Table 2, the Self-Regularization framework consistently outperforms its variants, “w/o Fine-tuning” and “w/o Const. Loss,” across all three datasets. This demonstrates that both the fine-tuning strategy and the constraint loss are crucial for effectively identifying unintended features and mitigating their influence on model predictions. Note that, the “w/o Const. Loss” variant even fails to significantly outperform the “w/o Regularization” baseline across the datasets. This supports our hypothesis in Section 3.3 that merely removing unintended features from the input space is insufficient, as their indirect effects on model predictions remain unaddressed.
Table 4. Performance of fine-tuned sparse autoencoders on Reward Modeling (RM), Toxic Detection (TD), and Disease Diagnosis (DD). We report the number of dead features (#Dead) and normalized mean square error (nMSE) in the representation reconstruction task. For both metrics, lower are better.
TD RM DD
#Dead nMSE #Dead nMSE #Dead nMSE
Pre-trained 29,49829{,}49829 , 498 0.4466 27,79627{,}79627 , 796 0.5049 33,98533{,}98533 , 985 0.5012
Fine-tuned 12,18512{,}18512 , 185 0.2288 15,86015{,}86015 , 860 0.3449 25,05625{,}05625 , 056 0.3337
Fine-tuning helps pre-trained SAE adapt to downstream tasks.
Table 4, fine-tuning significantly improves the performance of the pre-trained sparse autoencoders across all three datasets in both metrics. Specifically, fine-tuning reduces the number of dead features (#Dead) by 58.69%, 42.94%, and 26.27% for three datasets, respectively. Similarly, it decreases the normalized MSE for representation reconstruction by 48.77% on TD, 31.69% on RM, and 33.42% on DD. In addition, in Table 2, we could observe that on the Disease Diagnosis task, “w/o Fine-Tuning” drops the performance of our framework almost equivalent to the “Early Stopping”, indicating that the pre-trained SAE cannot provide fine-grained enough features for this task. These results highlight the importance of fine-tuning in adapting the pre-trained SAE to downstream tasks.
4.4. RQ3: Case Studies
4.4.1. Settings.
We provide some learned features from our fine-tuned sparse autoencoder for the Toxic Detection task in Table 5 to check the interpretability of the proposed framework. Since the toxic detection task aims to detect harmful intentions from user queries, we define the unintended features of this task as those features that are not semantically correlated to harmful behaviors. For each feature, we provide the most activated text spans on the training dataset, followed by the summary of the text spans and the relevance to harmful behaviors, labeled by GPT-4o-mini.
4.4.2. Results.
LLMs can identify unintended features according to the feature explanations.
We first focus on the Most Activated Text Spans in Table 5, and we could observe that the collected text spans for each feature typically refer to the same pattern. For example, the text spans for the first feature include various expressions of the concept “requesting someone’s home”. In addition, we could find that the LLM-generated summary of text spans and LLM-judged Safety Relevant look reasonable to humans. For instance, in the first feature, LLM successfully identifies its relevance to safety as the activated text spans indicate that the user is trying to locate someone’s house, which violates the privacy requirement. Similarly, we could observe intentions of non-violent crimes and terrorist attacks from the second and third features, and the LLM judges them as safety-relevant as our expectation. In contrast, the feature on the specific text span “According” and the feature on “repairing items” are not classified as safety-relevant. These observations emphasize the rationale of identifying unintended features by using LLMs.
4.5. RQ4: Sensitivity Analysis
4.5.1. Settings.
There are two critical hyper-parameters in our proposed framework, namely the Layer $l$ of extracted representations and $\beta$ in Equation (6). The sensitivity analysis on them is conducted using the reward modeling task, and we report the F1 score for different $l$ and $\beta$ in Figure 3 and Figure 4, respectively.
4.5.2. Results.
Layer=8Layer=16Layer=24Layer=32505050606060707070808080F1 Score (%)Early StoppingSelf-Regularization
Figure 3. Performance of regularized classifiers by using text embeddings generated from different layers of Mistral-7B-inst (32 layers in total) on the Reward Modeling Task.
Middle layers yield optimal performance due to richer semantic representations.
The performance across different layers in Figure 3 shows a reverse U-shape, with both the baseline and ours achieving their highest F1 scores at middle layers (e.g., Layer 16 and Layer 24). This trend reflects the richer and more task-relevant semantic information encoded in intermediate layers of the Mistral-7B model, aligned with findings from previous work (Wu et al., 2024). As the baseline achieves its peak performance at Layer 16, we conduct all experiments using embeddings from this layer.
[…]
Figure 4. Performance of regularized classifiers with different hyper-parameters $\beta$ on the Reward Modeling task.
Self-Regularization is robust to embeddings from different LLM layers $l$.
Figure 3 shows that classifiers trained with the proposed framework consistently outperform the baseline across all tested layers. Although the absolute F1 scores vary by layer, the relative improvement achieved by self-regularization remains significant, demonstrating the framework’s robustness to hidden representations derived from different layers.
Self-Regularization demonstrates robustness to the hyper-parameter $\beta$.
Figure 4 illustrates that the proposed method consistently achieves an F1 score of at least 70.0% on the Reward Modeling Task across a wide range of $\beta$ values, significantly outperforming the baseline “Early STopping” at 69.01%. Notably, when $\beta$ is set between 0.75 and 6.0, the method maintains peak performance, with the highest F1 score of approximately 72.44% achieved at $\beta=3.0$. This observation indicates that the proposed framework is stable and not overly sensitive to the choice of $\beta$.
5. Related Work
Our work builds on research in LLM interpretability and classifier regularization. Efforts to explain LLMs have leveraged activation monitoring (Wang et al., 2023a; Jin et al., 2024, 2025, 2024), probing (Belinkov et al., 2018; Jawahar et al., 2019), and basis decomposition (Elhage et al., 2022; Olah et al., 2020) to address neuron polysemanticity (Arora et al., 2018). Recent advances use sparse autoencoders (Bricken et al., 2023; Cunningham et al., 2023) to extract monosemantic features, improving interpretability and enabling control over model behaviors, with applications extending to translation (Dumas et al., [n. d.]), circuit detection (Marks et al., 2024), model steering (Wu et al., 2025; Shu et al., 2025a; Zhao et al., 2025), and scaling to larger LLMs (Templeton et al., 2024; Gao et al., 2024; Lieberum et al., 2024). We refer readers to a survey on SAEs (Shu et al., 2025b) for the latest progress of studying SAEs. We extend this line by fine-tuning sparse autoencoders on domain data to mitigate dead neurons and enhance task-specific feature extraction. In classifier regularization, traditional methods (Hoerl and Kennard, 1970; Srivastava et al., 2014) prevent model overfitting, while modern approaches such as spectral normalization (Yoshida and Miyato, 2017) and focal loss (Ross and Dollár, 2017) refine model generalization. Our work differs by leveraging sparse autoencoders to achieve semantic regularization on embedding-based classifier training.
6. Conclusion
This paper introduces a framework that regularizes the usage of unintended features for embedding-based text classifiers. Our approach identifies unintended features within LLM-generated embeddings with sparse autoencoders, and regularizes their impact for classification by introducing a constraint loss. Empirical evaluations across three challenging text classification tasks demonstrate that our method effectively reduces unintended feature reliance and improves the generalizability of classifiers. This work steps a solid stamp toward controllable embedding-based text classification.
Appendix A Minimum Sample Size to Effectively Estimate Mean of Sparse Feature Activations
In this section, we aim to quantify the minimum number of training samples required to obtain a reliable estimate of the average activation values $\mathbf{A}=\sigma(\mathbf{X}\cdot\mathbf{W})$ of learned sparse features from sparse autoencoders. With the sparse nature of these learned features, we assume that each learned sparse feature is activated with probability $p$ within an input text $x$, and when activated, the value falls a normal distribution $\mathcal{N}(\mu,\sigma^{2})$, the estimation of $\mu$ requires a sufficient number of nonzero activations $n_{\text{sparse}}$. Formally, the activations of sparse features follow a Bernoulli-Gaussian distribution:
(7) $\mathbf{A}=\begin{cases}Z,&\text{with probability }p,\\ 0,&\text{with probability }1-p,\end{cases}$
where $Z\sim\mathcal{N}(\mu,\sigma^{2})$. The key challenge in estimating $\mu$ arises from the sparsity induced by $p$, which effectively reduces the number of informative samples (the values are greater than 0).
To determine the minimum required training samples $n_{\text{sparse}}$, we recall that the effective sample size $n_{\text{normal}}$ to estimate the mean of a normal distribution with a margin of error $d$ at confidence level $1-\alpha$ is given by the Central Limit Theorem (Fischer, 2011) as:
(8) $n_{\text{normal}}=\left(\frac{z_{\alpha/2}\cdot\sigma}{d}\right)^{2},$
where $z_{\alpha/2}$ is the z-score corresponding to the significance level $\alpha$ under a normal distribution. However, since only a fraction $p$ of the total samples yield nonzero activations, the expected number of usable samples is $p\cdot n_{\text{normal}}$. Therefore, the sample size for estimating the mean activations of sparse features is adjusted as:
(9) $n_{\text{sparse}}=\frac{1}{p}\left(\frac{z_{\alpha/2}\cdot\sigma}{d}\right)^{2}.$
[…]
(10) $n_{\text{sparse}}=\frac{1}{0.01}\left(\frac{1.96\times\sigma}{0.1\times\sigma}\right)^{2}\approx 38416.$
Given that real-world datasets (Wang, 2018) usually consist of thousands of training samples, which is significantly smaller than $n_{\text{sparse}}\approx 38416$, effectively estimating the average activations $\mathbf{A}$ of the learned sparse feature is infeasible in practice.
Table 5. Additional qualitative analysis on LLM-identified intended/unintended features for the Reward Modeling task, where the classifier predicts the helpfulness of a chatbot’s response to the user. We highlight phrases in the text spans that are semantically correlated to helpful behaviors, such as “being polite”, “considering diverse cultures”, and “solving math problems”.
Most Activated Text Spans Summary of Text Spans Helpful Relevant
Span 1: _[BOT]:__ It could; Span 2: __ Is it possible that your symptoms could be related; Span 4: __[BOT]:__ Perhaps there’s; Span 4: BOT]:__ Good question! It might be; Span 5: [BOT]:__ Do you think it might; Span 6: s a possibility of damage, but it could also; Span 7: BOT]:__ If you are, you might; Span 8: n__[BOT]:__ Maybe you’; Span 9: n__[BOT]:__ It could be; Span 10: __[BOT]:__ Could it Expressing a possibility or uncertainty. Yes
Span 1: a Japanese mayonnaise called “Nozaw; Span 2: is popularly served with Katogo (beef; Span 3: ame, Spyro: Year of the Dragon; Span 4: upgrade pack called Diablo II: Lord of Dest; Span 5: two hours after giving the infant iron-containing; Span 6: also listen to George Harrison’s ”My Sweet; Span 7: ina is Tina Turner’s autobiography; Span 8: model name, for example the 901; Span 9: I suggest looking for pieces called Goldberg Vari; Span 10: a message like “The bear ate the gummy The provided text Spans contain references to diverse cultural items (Japanese mayonnaise), food (Katogo), and video games (Spyro: Year of the Dragon, Diablo II: Lord of Destruction). Yes
Span 1: day will be 24 hours, 1; Span 2: 20 years, you will have \$1; Span 3: 37 degrees Fahrenheit = 3; Span 4: Sure. 3.5 rounded up is ; Span 5: 37 degrees Fahrenheit = 3 The feature captures numerical values and conversions. Yes
Span 1: Was considering; Span 2: I’m thinking of getting; Span 3: I was thinking about getting; Span 4: I’m thinking about getting; Span 5: What should I look for; Span 6: Why are there; Span 7: What things should I look for; Span 8: Should I get; Span 9: I’m trying to decide whether; Span 10: How do I decide between getting The feature captures user general queries. No
Appendix B Details of Training Sparse Autoencoders
B.1. Pre-training Sparse Autoencoders.
[…]
B.2. Fine-tuning Sparse Autoencoders.
[…]
Appendix C Details of Identifying Unintended Features
C.1. Interpreting Learned Features
Building on prior work in LLM-as-a-judge (Bills et al., 2023; Chaudhary and Geiger, 2024; Gao et al., 2024; Lieberum et al., 2024), we interpret the learned feature vectors from fine-tuned sparse autoencoders by identifying the Top-10 text spans that most activate each feature, with each span limited to a maximum of 32 tokens. To summarize the underlying patterns of these activations, we employ gpt-4o-mini-2024-07-18 as our machine annotator, with a temperature of 0 for deterministic decoding. Each generated response is capped at 1024 tokens. To enhance the reliability of this process, we design a structured prompting framework that incorporates a role-playing strategy and in-context examples. Following previous work (Bills et al., 2023), our machine annotator has the option to respond with “Cannot Tell” if it detects no meaningful patterns among the activated text spans. Additionally, to mitigate hallucinations, we prompt the LLM in a separate thread to verify whether its previously generated summary accurately reflects the underlying data.
C.2. Identifying Unintended Features
For all three tasks, we define unintended features as those that lack a clear semantic correlation with the task, as determined by human-crafted evaluation rubrics. Specifically, we reference prior work to establish these rubrics: toxic detection follows (Llama Team, 2024), reward modeling is based on (Ouyang et al., 2022), and disease diagnosis adheres to guidelines from (Xu et al., 2019). To assess feature-task relevance, we adopt the annotation framework from (Lieberum et al., 2024), where the machine annotator classifies correlations into four levels: “Yes”, “Probably”, “Maybe”, and “No”. Features are considered unintended if they are are assigned a task relevance level lower than “Probably”.
Appendix D Extended Case Study on Reward Modeling
Table 5 provides an extended qualitative analysis of LLM-identified intended and unintended features for the Reward Modeling task. In this task, the classifier predicts the helpfulness of a chatbot’s responses to users. Unintended features are defined as those that are not semantically aligned with the goal of classifying helpfulness. A particularly noteworthy example is the second feature, where the raw explanations highlight a diverse range of cultural references.Conversely, other features such as the fourth example (focused on repetitive phrases like “over the”) are identified as unintended, as they lack meaningful alignment with the task.